OpenStudy (anonymous):

a typist types 900 words in 60 minutes.What is her rate of typing per minutes?

OpenStudy (anonymous):

how do u get 1 minute from 60 minutes?

OpenStudy (arnavguddu):

in 60 min typist types 900 words in 1 min .... types 900/60 = 15 words so 15 words per min

OpenStudy (hoblos):

900/60 = 15 words / min

OpenStudy (anonymous):

900/60 = 15 words / min :D

OpenStudy (anonymous):

w8 i twll u 60 mins = 900 words 1 min=900/60 = 15 words

OpenStudy (anonymous):

yes that is right @rosabelle